OpenStudy (anonymous):

Simplify : (4) / (6-2 sqrt6) A. 1+ (sqrt6 i) / (3) B. 3+ (sqrt3 i) / (3) C. 2+ (3 sqrt6 i) / (4) D. 2- (sqrt6) / (6) E. none of these I'm getting 2+ (2 sqrt6) / (3) so E. none of these?

OpenStudy (anonymous):

\[\frac{ 4 }{6-2\sqrt{6} }\]

OpenStudy (anonymous):

A. \[1+ \frac{ \sqrt{6}i }{ 3 }\] B. \[3+\frac{ \sqrt{3}i }{ 3 }\] C. \[2+\frac{ 3\sqrt{6}i }{ 4 }\] D. \[2-\frac{ \sqrt{6} }{ 6 }\]

OpenStudy (anonymous):

i'm getting \[2+ \frac{ 2\sqrt{6} }{ 3 }\]

OpenStudy (alyssa_xo):

How did you get that?

OpenStudy (anonymous):

hold on that didnt show up well. let me do it under equation.

OpenStudy (anonymous):

4(6+2 sqrt6)/(6-2 sqrt6)(6+2 sqrt6) (24+8 sqrt6)/(36-24) (24+8 sqrt6)/12 2+ (2 sqrt6) / (3)

OpenStudy (anonymous):

there we go.

OpenStudy (anonymous):

@alyssa_xo

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

@jigglypuff314